Use matrices and row operations to solve the following system of equations:

2x-y+3z=7

x-y-z=0

-3x-2z=-11
Answer #1

for matrix A =

2 -1 3
1 -1 -1
-3 0 -2

and matrix B =

x
y
z

also C

7
0
-11

here AB =C

therefore B =A-1C

for inverse of A:

-0.2 0.2 -0.4
-0.5 -0.5 -0.5
0.3 -0.3 0.1

therefore multiplying above with C:

A-1C =

3
2
1

hence x =3

y=2

z=1